OpenStudy (anonymous):

@ali2x2 @robtobey @jigglypuff314 @heretohelpalways The equation of the line of best fit of a scatter plot is y = –3x + 2. What is the slope of the equation? –3 –2 2 3

jigglypuff314 (jigglypuff314):

They gave you the equation in slope intercept form :) where \( y = mx + b \) and \(m\) is the slope

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so what do you think it is @kool_22?

OpenStudy (anonymous):

2

OpenStudy (sleepyhead314):

not quite :) compare y = -3x + 2 to y = mx + b ^ what does "m" line up with?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

The correct answer is -3.

OpenStudy (sleepyhead314):

that's right :)